OpenStudy (anonymous):

Solve the equation -3(x-5)=-5(-3-4x) a. 0 b. -3 c. all real numbers d. no solution

OpenStudy (anonymous):

plz help me

OpenStudy (abb0t):

Start by distributing the 3 across the parenthesis. So you have: \[(-3 \times x) + (-3 \times 5) = (-5 \times -3) + (-5 \times -4x)\]

OpenStudy (abb0t):

Do the same with both sides. Which is what I did there.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

whats the answer

OpenStudy (abb0t):

That's where we're going with this. Did you distribute? After you distribute, you have -3x+15 = 15+20x Your next step is to use your algebra rules to massage this a bit. So basically, subtract 15 from both sides, and you get -3x = 20x Now, notice how you can divide both sides to get a constant value w/ no x = term? So what do you think the answer might be?